# Closed-Form Feedback-Free Learning with Forward Projection

### v3

@@ A.6 Computing the Forward Projection estimator
Neural network models often manage storage demands by utilizing data mini-batches, allowing them to train on arbitrarily large datasets without a corresponding increase in memory requirements. Although the Forward Projection estimator described in equation (2) includes large matrix terms $\textbf{A}\in\mathbb{R}^{n\times m_{l-1}}$ and $\textbf{Z}\in\mathbb{R}^{n\times m_{l-1}}$, these do not need to be loaded into memory. Assume $\textbf{A}_{l-1}\in\mathbb{R}^{N\times m_{l-1}}$ and $\textbf{Y}\in\mathbb{R}^{N\times m_{L}}$ are available as a set of row-vectors $\{\textbf{a}_{i,l-1}\}^{N}_{i=1}\subset\mathbb{R}^{1\times m_{l-1}}$. and $\{\textbf{y}_{i}\}^{N}_{i=1}\subset\mathbb{R}^{1\times m_{L}}$. The matrix cross product is the sum of the row outer products, such that
$\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}=\sum^{N}_{i=1}\textbf{a}_{i,l-1}\textbf{a}_{i,l-1}^{T}.$ (22)
Therefore, the $\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}\in\mathbb{R}^{m_{l-1}\times m_{l-1}}$ term may be accumulated sequentially over instances $\{\textbf{a}_{1,l-1},\ldots,\textbf{a}_{N,l-1}\}$, with $\mathcal{O}(m_{l-1}^{2})$ memory and $\mathcal{O}(Nm_{l-1}^{2})$ compute, such that
$(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1})^{(i)}=(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1})^{(i-1)}+\textbf{a}_{i,l-1}\textbf{a}_{i,l-1}^{T},$ (23)
where $(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1})^{(0)}$ is zero matrix of dimension $m_{l-1}\times m_{l-1}$. Assume $\textbf{Y}\in\mathbb{R}^{N\times m_{L}}$ is available as $\{\textbf{y}_{i}\}^{N}_{i=1}\subset\mathbb{R}^{1\times m_{L}}$. We may generate the corresponding target potentials $\{\tilde{\textbf{z}}_{i,l-1}\}^{N}_{i=1}\subset\mathbb{R}^{1\times m_{l}}$, using predetermined matrices $\textbf{Q}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$ and $\textbf{U}_{l}\in\mathbb{R}^{m_{L}\times m_{l}}$according to (4). Thus, $\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$ may also be accumulated sequentially such that
$\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}_{l}=\sum^{N}_{i=1}\textbf{a}_{i,l-1}\tilde{\textbf{z}}_{i,l-1}^{T},$ (24)
requiring $\mathcal{O}(m_{l-1}m_{l})$ memory and $\mathcal{O}(Nm_{l-1}m_{l})$ compute. Following sequential computation of equations (23) and (24), equation (25) may be computed with one order $\mathcal{O}(m_{l-1}^{3})$ matrix inversion and one order $\mathcal{O}(m_{l-1}^{2}m_{l})$ matrix multiplication, yielding:
$\textbf{W}_{l}=(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}+\lambda\textbf{I})^{-1}(\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}_{l}).$ (25)
@@ A.7 Information Encoding and Interpretability
We now consider the information encoding properties of Forward Projection training. Assuming $g_{l}^{-1}$ exists everywhere, equation (4) relates labels $\textbf{y}\in\mathbb{R}^{1\times m_{L}}$ to target potentials $\tilde{\textbf{z}}_{l}\in\mathbb{R}^{1\times m_{l}}$, given presynaptic inputs $\textbf{a}_{l-1}\in\mathbb{R}^{1\times m_{l-1}}$ via projection matrices $\textbf{Q}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$ and $\textbf{U}_{l}\in\mathbb{R}^{m_{L}\times m_{l}}$, such that
$\textbf{y}\textbf{U}_{l}=g_{l}^{-1}\left(\tilde{\textbf{z}}_{l}-g_{l}(\textbf{a}_{l-1}\textbf{Q}_{l})\right).$ (26)
Letting $\textbf{U}_{l}^{+}$ denote the Moore-Penrose inverse, and assuming that $\textbf{U}_{l}\textbf{U}_{l}^{+}\approx I$, we may reconstruct y, such that
$\textbf{y}\approx g_{l}^{-1}\left(\tilde{\textbf{z}}_{l}-g_{l}(\textbf{a}_{l-1}\textbf{Q}_{l})\right)\textbf{U}_{l}^{+}.$ (27)
Assuming realised neuron pre-activation potentials $\textbf{z}_{l}$ are a good approximation of the target potentials, i.e., $\textbf{z}_{l}\approx\tilde{\textbf{z}}_{l}$, we may estimate the label from the hidden layer, such that
$\hat{\textbf{y}}_{l}:=g_{l}^{-1}\left(\textbf{z}_{l}-g_{l}(\textbf{a}_{l-1}\textbf{Q}_{l})\right)\textbf{U}_{l}^{+}.$ (28)
Thus, neural potentials are interpretable layer-wise as label predictions. Likewise, pre-synaptic inputs may be estimated such that
$\hat{\textbf{a}}_{l-1}:=g_{l}^{-1}\left(\textbf{z}_{l}-g_{l}(\textbf{y}\textbf{U}_{l})\right)\textbf{Q}_{l}^{+}.$ (29)
@@
The function $g_{l}:=\operatorname{sign}(x)$ was selected due to computational simplicity and predictable distribution. The $\operatorname{tanh}$ function was employed as a surrogate approximation for $g_{l}^{-1}$, yielding satisfactory results.
@@ A.8 Stability Experiments
Figure 6: A: Test accuracy of hidden layer explanations in a $4\times 1000$ neuron MLP trained by FP on FMNIST. B: Test accuracy of $3\times 1000$ neuron MLP trained trained by FP on FMNIST using different $\lambda$ regularisation values. C: Test accuracy of $3\times 1000$ neuron MLP trained trained by FP on FMNIST with additive Gaussian noise in training data. D: Variability of FP training performance using different random projections in $3\times 1000$ neuron MLP on FMNIST. Points represent mean over all replicates and error bars represent $\pm$ one standard deviation. BP: backpropagation; DTP: Difference Target Propagation; $F$: Forward-Forward; FP: Forward Projection; LP: label projection; LPN: noisy label projection; LS: Local Supervision; PC: Predicitve Coding; RF: random features.
@@ A.9 Optimisation Theory
We now compare the fit of the Forward Projection estimator in a two-layer network, to that of a simple linear estimator in random data. Consider a dataset $\{\textbf{x}_{i},\textbf{y}_{i}\}_{i=1}^{N}$ drawn from a multivariate Gaussian distribution, where $\textbf{x}_{i}\sim\mathcal{N}(\textbf{0},\textbf{I}_{m})$ and the independent target $\textbf{y}_{i}\sim\mathcal{N}(\textbf{0},\textbf{I}_{d})$. Assume $N>m>d$. We consider a fully connected neural network with a single hidden layer of dimension $m$, with ReLU activation. The task is to identify network weights $\{\textbf{W}_{1},\textbf{W}_{2}\}$ to minimise the error $\lVert\textbf{Y}-\operatorname{ReLU}(\textbf{XW}_{1})\textbf{W}_{2}\rVert$. We first consider the expected error of a single layer, linear model. For convenience, we will set the penalty factor $\lambda=0$, such that
$\textbf{W}_{1}:=(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}\textbf{Y}.$ (31)
We denote the projection matrix of X as
$\textbf{P}_{\textbf{x}}=\textbf{X}(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}.$ (32)
@@
Therefore, the spectrum of $\textbf{P}_{\textbf{x}}$ is composed of a unit eigenvalue with multiplicity $m$ and a zero eigenvalue with multiplicity NmN-m. Letting $\hat{\textbf{Y}}_{1}=\textbf{P}_{\textbf{x}}\textbf{Y}$, we measure the error $\textbf{E}_{1}=\textbf{Y}-\textbf{P}_{\textbf{x}}\textbf{Y}$, such that
$\lVert\textbf{E}_{1}\rVert=\lVert(\textbf{I}-\textbf{P}_{\textbf{x}})\textbf{Y}\rVert.$ (33)
The idempotent matrix $(\textbf{I}-\textbf{P}_{\textbf{x}})$ has rank NmN-m, as it is a projection onto the orthogonal complement of the column space of X. Therefore, the expected squared Frobenius norm of the error is given by
@@
$\mathbb{E}\left[\frac{\lVert\textbf{E}_{1}\rVert^{2}_{F}}{\lVert\textbf{Y}\rVert^{2}_{F}}\right]=\frac{N-m}{N}.$ (34)
Let $\textbf{U}\in\mathbb{R}^{d\times m}$ and $\textbf{Q}\in\mathbb{R}^{m\times m}$ . Letting $\tilde{\textbf{Z}}=\textbf{XQ}+\textbf{YU}$, we now fit $\textbf{W}_{1}$ to estimate $\tilde{\textbf{Z}}$ from X, such that
$\textbf{W}_{1}:=(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}(\textbf{XQ}+\textbf{YU}).$ (35)
As $(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}\textbf{X}=\textbf{I}$, we have
$\textbf{W}_{1}=\textbf{Q}+(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}\textbf{YU}.$ (36)
Thus, the pre-activation potential at the first layer is given by:
$\textbf{X}\textbf{W}_{1}=\textbf{XQ}+\textbf{P}_{\textbf{x}}\textbf{YU}.$ (37)
Assuming U has full row rank, we have $\textbf{UU}^{+}=\textbf{I}_{d}$, as $m>d$. Furthermore, as Q and U have random Gaussian entries, we have
$\mathbb{E}[\textbf{X}\textbf{W}_{1}\textbf{U}^{+}]=\mathbb{E}[\textbf{XQU}^{+}]+\mathbb{E}[\textbf{P}_{\textbf{x}}\textbf{YUU}^{+}]=\textbf{P}_{\textbf{x}}\textbf{Y}=\hat{\textbf{Y}}_{1}.$ (38)
We now consider the covariance $\operatorname{Cov}(\textbf{XW}_{1},\textbf{Y})$. We assume $\textbf{XW}_{1}$ has zero mean. We have
$\mathbb{E}[\langle\textbf{XW}_{1},\textbf{Y}\rangle]=\mathbb{E}[\operatorname{tr}\left((\textbf{XQ}+\textbf{P}_{\textbf{x}}\textbf{YU})^{T}\textbf{Y}\right)]$ (39)
By independence of Q, we have $\mathbb{E}[\operatorname{tr}((\textbf{XQ})^{T}\textbf{Y})]=0$. The second term expands to
$\operatorname{tr}\left((\textbf{P}_{\textbf{x}}\textbf{YU})^{T}\textbf{Y}\right)=\operatorname{tr}\left(\textbf{U}^{T}\textbf{Y}^{T}\textbf{P}_{\textbf{x}}\textbf{Y})\right)$ (40)
As $\textbf{P}_{\textbf{x}}$ is a projector matrix with rank $m$, and Y is Gaussian, we have
$\mathbb{E}[\textbf{Y}^{T}\textbf{P}_{\textbf{x}}\textbf{Y}]=d\cdot\operatorname{tr}(\textbf{P}_{\textbf{x}})\textbf{I}_{d}=dm\cdot\textbf{I}_{d}$ (41)
@@
$\mathbb{E}[\langle\textbf{XW}_{1},\textbf{Y}\rangle]=\operatorname{tr}\left(\textbf{U}^{T}\textbf{Y}^{T}\textbf{P}_{\textbf{x}}\textbf{Y})\right)=dm\cdot\operatorname{tr}(\textbf{U}^{T})$ (42)
As $m>d$ and U has random standard Gaussian entries, we have $\operatorname{Var}[\operatorname{tr}(\textbf{U}^{T})]=d$. Therefore, we have
$\mathbb{E}[\lvert\langle\textbf{XW}_{1},\textbf{Y}\rangle\rvert]=\lvert\operatorname{tr}\left(\textbf{U}^{T}\textbf{Y}^{T}\textbf{P}_{\textbf{x}}\textbf{Y})\right)\rvert\sim dm\sqrt{d}.$ (43)
After normalising for $\lVert\textbf{XW}_{1}\rVert_{F}\lVert\textbf{Y}\rVert_{F}\sim\sqrt{Nm}\cdot\sqrt{Nd}$, the correlation magnitude scales according to
$\mathbb{E}\left[\lvert\operatorname{Corr}(\textbf{XW}_{1},\textbf{Y})\rvert\right]\sim\frac{\operatorname{tr}\langle\textbf{XW}_{1},\textbf{Y}\rangle}{\lVert\textbf{XW}_{1}\rVert_{F}\lVert\textbf{Y}\rVert_{F}}=\frac{d\sqrt{m}}{N}.$ (44)
Assuming $\operatorname{rank}(\textbf{A})=m$, we may define $\textbf{P}_{\textbf{a}}=\textbf{A}(\textbf{A}^{T}\textbf{A})^{-1}\textbf{A}^{T}$. We note that A is dependent on Y with a non-linear relationship, which is induced by the $\operatorname{ReLU}$ activations. We now consider modelling Y in $\mathcal{C}(\textbf{A})\cap\mathcal{C}(\textbf{X})^{\perp}$, the subspace of A orthogonal to X. For large $N$, application of $\operatorname{ReLU}$ nonlinearities in the construction of A may yield $\operatorname{dim}\left(\mathcal{C}(\textbf{A})\cap\mathcal{C}(\textbf{X})^{\perp}\right)\approx m$. Recall from (44) that the column space of $\textbf{XW}_{1}$ is correlated with Y. As $\textbf{A}=\operatorname{ReLU}(\textbf{XW}_{1})$, the column space of A remains correlated with Y. Therefore

### v4

@@ A.6 Computing the Forward Projection estimator
Neural network models often manage storage demands by utilizing data mini-batches, allowing them to train on arbitrarily large datasets without a corresponding increase in memory requirements. Although the Forward Projection estimator described in equation (2) includes large matrix terms $\mathbf{A}_{l-1}\in\mathbb{R}^{N\times m_{l-1}}$ and $\tilde{\mathbf{Z}}_{l}\in\mathbb{R}^{N\times m_{l}}$, these do not need to be loaded into memory. Assume $\mathbf{A}_{l-1}\in\mathbb{R}^{N\times m_{l-1}}$ and $\mathbf{Y}\in\mathbb{R}^{N\times m_{L}}$ are available as a set of row-vectors $\{\mathbf{a}_{i,l-1}\}^{N}_{i=1}\subset\mathbb{R}^{1\times m_{l-1}}$ and $\{\mathbf{y}_{i}\}^{N}_{i=1}\subset\mathbb{R}^{1\times m_{L}}$. The Gramian matrix is the sum of the row outer products, such that
$\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1}=\sum^{N}_{i=1}\mathbf{a}_{i,l-1}^{\top}\mathbf{a}_{i,l-1}.$ (23)
Therefore, the $\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1}\in\mathbb{R}^{m_{l-1}\times m_{l-1}}$ term may be accumulated sequentially over instances $\{\mathbf{a}_{1,l-1},\ldots,\mathbf{a}_{N,l-1}\}$, with $\mathcal{O}(m_{l-1}^{2})$ memory and $\mathcal{O}(Nm_{l-1}^{2})$ compute, such that
$(\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1})^{(i)}=(\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1})^{(i-1)}+\mathbf{a}_{i,l-1}^{\top}\mathbf{a}_{i,l-1},$ (24)
where $(\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1})^{(0)}$ is the zero matrix of dimension $m_{l-1}\times m_{l-1}$. We may generate the corresponding target potentials $\{\tilde{\mathbf{z}}_{i,l}\}^{N}_{i=1}\subset\mathbb{R}^{1\times m_{l}}$, using predetermined matrices $\mathbf{Q}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$ and $\mathbf{U}_{l}\in\mathbb{R}^{m_{L}\times m_{l}}$according to (4). Thus, $\mathbf{A}_{l-1}^{\top}\tilde{\mathbf{Z}}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$ may also be accumulated sequentially such that
$\mathbf{A}_{l-1}^{\top}\tilde{\mathbf{Z}}_{l}=\sum^{N}_{i=1}\mathbf{a}_{i,l-1}^{\top}\tilde{\mathbf{z}}_{i,l},$ (25)
requiring $\mathcal{O}(m_{l-1}m_{l})$ memory and $\mathcal{O}(Nm_{l-1}m_{l})$ compute. Addition of the $\lambda\mathbf{I}$ regularisation term has negligible $\mathcal{O}(m_{l-1})$ overhead. Following sequential computation of equations (24) and (25), equation (2) may be computed with one order $\mathcal{O}(m_{l-1}^{3})$ matrix inversion and one order $\mathcal{O}(m_{l-1}^{2}m_{l})$ matrix multiplication, yielding:
$\mathbf{W}_{l}=(\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1}+\lambda\mathbf{I})^{-1}(\mathbf{A}_{l-1}^{\top}\tilde{\mathbf{Z}}_{l}).$ (26)
@@ A.7 Information Encoding and Interpretability
We now consider the information encoding properties of Forward Projection training. Assuming $g_{l}^{-1}$ exists everywhere, equation (4) relates labels $\mathbf{y}\in\mathbb{R}^{1\times m_{L}}$ to target potentials $\tilde{\mathbf{z}}_{l}\in\mathbb{R}^{1\times m_{l}}$, given presynaptic inputs $\mathbf{a}_{l-1}\in\mathbb{R}^{1\times m_{l-1}}$ via projection matrices $\mathbf{Q}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$ and $\mathbf{U}_{l}\in\mathbb{R}^{m_{L}\times m_{l}}$, such that
$\mathbf{y}\mathbf{U}_{l}=g_{l}^{-1}\left(\tilde{\mathbf{z}}_{l}-g_{l}(\mathbf{a}_{l-1}\mathbf{Q}_{l})\right).$ (27)
Letting $\mathbf{U}_{l}^{+}$ denote the Moore-Penrose inverse, and assuming that $\mathbf{U}_{l}\mathbf{U}_{l}^{+}=\mathbf{I}$, we may reconstruct $\mathbf{y}$, such that
$\mathbf{y}=g_{l}^{-1}\left(\tilde{\mathbf{z}}_{l}-g_{l}(\mathbf{a}_{l-1}\mathbf{Q}_{l})\right)\mathbf{U}_{l}^{+}.$ (28)
Assuming realised neuron pre-activation potentials $\mathbf{z}_{l}$ are a good approximation of the target potentials, i.e., $\mathbf{z}_{l}\approx\tilde{\mathbf{z}}_{l}$, we may estimate the label from the hidden layer, such that
$\hat{\mathbf{y}}_{l}:=g_{l}^{-1}\left(\mathbf{z}_{l}-g_{l}(\mathbf{a}_{l-1}\mathbf{Q}_{l})\right)\mathbf{U}_{l}^{+}.$ (29)
Thus, neural potentials are interpretable layer-wise as label predictions. Assuming $\mathbf{Q}_{l}\mathbf{Q}_{l}^{+}\approx\mathbf{I}$, pre-synaptic inputs may also be estimated approximately such that
$\hat{\mathbf{a}}_{l-1}:=g_{l}^{-1}\left(\mathbf{z}_{l}-g_{l}(\mathbf{y}\mathbf{U}_{l})\right)\mathbf{Q}_{l}^{+}.$ (30)
@@
The element-wise function $g_{l}:=\operatorname{sign}(x)$ was selected due to computational simplicity and predictable distribution. The $\operatorname{tanh}$ function was employed as a heuristic surrogate for $g_{l}^{-1}$, yielding satisfactory results.
@@ A.8 Stability Experiments
Figure 6: Technical analysis of multilayer perceptron trained by Forward Projection. A: Test accuracy of hidden layer explanations in a $4\times 1000$ neuron MLP trained by FP on FMNIST. B: Test accuracy of $3\times 1000$ neuron MLP trained trained by FP on FMNIST using different $\lambda$ regularisation values. C: Test accuracy of $3\times 1000$ neuron MLP trained trained by FP on FMNIST with additive Gaussian noise in training data. D: Variability of FP training performance using different random projections in $3\times 1000$ neuron MLP on FMNIST. Points represent mean over all replicates and error bars represent $\pm$ one standard deviation. BP: backpropagation; DTP: Difference Target Propagation; FF: Forward-Forward; FP: Forward Projection; LP: label projection; LPN: noisy label projection; LS: Local Supervision; PC: Predicitve Coding; RF: Random Features.
@@ A.9 Optimisation Theory
We now compare the ordinary least squares linear model to a two-layer network trained by Forward Projection and a Random Features model. We aim to demonstrate analytically that the expected error of the Forward Projection model improves upon the others.
Consider a dataset $\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$, where $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{m_{l}})$ and target $\mathbf{y}_{i}\in\mathbb{R}^{m_{L}}$. Let $\mathbf{X}\in\mathbb{R}^{N\times m_{l}}$ denote the design matrix with rows $\mathbf{x}_{i}^{\top}$ and $\mathbf{Y}\in\mathbb{R}^{N\times m_{L}}$ denote the target matrix with rows $\mathbf{y}_{i}^{\top}$. Let $\operatorname{col}(\mathbf{X})$ denote the column space of matrix $\mathbf{X}$, the span of its column vectors. We assume $N>m_{l}\gg m_{L}$ and that $\mathbf{X}$ has full column rank. We first consider the expected error of a single-layer, linear model, with weights $\hat{\mathbf{W}}_{\mathrm{linear}}\in\mathbb{R}^{m_{l}\times m_{L}}$. We estimate $\hat{\mathbf{W}}_{\mathrm{linear}}$, by ordinary least squares [46], such that
$\hat{\mathbf{W}}_{\mathrm{linear}}:=(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}\mathbf{Y}.$ (32)
We denote the projection matrix of $\mathbf{X}$ as
$\mathbf{P}_{\mathbf{X}}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}.$ (33)
@@
The spectrum of the idempotent projection matrix $\mathbf{P}_{\mathbf{X}}$ is composed of a unit eigenvalue with multiplicity $m_{l}$ and a zero eigenvalue with multiplicity $N-m_{l}$ [47]. Letting $\hat{\mathbf{Y}}_{\mathrm{linear}}=\mathbf{P}_{\mathbf{X}}\mathbf{Y}$, the model error is given by
@@
$E_{\mathrm{linear}}=\mathbb{E}\left[\lVert\mathbf{Y}-\hat{\mathbf{Y}}_{\mathrm{linear}}\rVert^{2}_{F}\right]=\mathbb{E}\left[\lVert(\mathbf{I}-\mathbf{P}_{\mathbf{X}})\mathbf{Y}\rVert^{2}_{F}\right],$ (34)
Letting $\mathbf{Y}_{\perp}=(\mathbf{I}-\mathbf{P}_{\mathbf{X}})\mathbf{Y}$, we decompose [48]
$\mathbf{Y}=\mathbf{P}_{\mathbf{X}}\mathbf{Y}+(\mathbf{I}-\mathbf{P}_{\mathbf{X}})\mathbf{Y}=\hat{\mathbf{Y}}_{\mathrm{linear}}+\mathbf{Y}_{\perp},$ (35)
where $\operatorname{col}(\hat{\mathbf{Y}}_{\mathrm{linear}})\subset\operatorname{col}(\mathbf{X})$, and $\mathbf{Y}_{\perp}$ contains residuals. Let $\operatorname{col}(\mathbf{X})^{\perp}$ denote the $(N-m_{l})$-dimensional space orthogonal to $\operatorname{col}(\mathbf{X})$. We note that as $(\mathbf{I}-\mathbf{P}_{\mathbf{X}})$ projects onto $\operatorname{col}(\mathbf{X})^{\perp}$, $\mathbf{Y}_{\perp}$ lies in $\operatorname{col}(\mathbf{X})^{\perp}$. By symmetric idempotency of $(\mathbf{I}-\mathbf{P}_{\mathbf{X}})$ and cyclic trace invariance [48], we have
$E_{\mathrm{Linear}}=\mathbb{E}\big[\lVert\mathbf{Y}_{\perp}\rVert_{F}^{2}\big]=\mathbb{E}\left[\operatorname{Tr}\left((\mathbf{I}-\mathbf{P}_{\mathbf{X}})\,\mathbf{Y}\mathbf{Y}^{\top}(\mathbf{I}-\mathbf{P}_{\mathbf{X}})^{\top}\right)\right]\\ =\mathbb{E}\left[\operatorname{Tr}\left((\mathbf{I}-\mathbf{P}_{\mathbf{X}})^{\top}(\mathbf{I}-\mathbf{P}_{\mathbf{X}})\,\mathbf{Y}\mathbf{Y}^{\top}\right)\right]=\mathbb{E}\left[\operatorname{Tr}\left((\mathbf{I}-\mathbf{P}_{\mathbf{X}})\,\mathbf{Y}\mathbf{Y}^{\top}\right)\right].$ (36)
We now consider a two-layer neural network model trained by forward projection for fitting the same data. The model has $m_{l}$ hidden $\operatorname{ReLU}$-activated neurons, such that
$\hat{\mathbf{Y}}_{\mathrm{FP}}:=\operatorname{ReLU}(\mathbf{X}\hat{\mathbf{W}}_{1})\hat{\mathbf{W}}_{2},$ (37)
The task is to identify weight matrices $\hat{\mathbf{W}}_{1}\in\mathbb{R}^{m_{l}\times m_{l}},\hat{\mathbf{W}}_{2}\in\mathbb{R}^{m_{l}\times m_{L}}$ to minimise the error $\lVert\mathbf{Y}-\hat{\mathbf{Y}}_{\mathrm{FP}}\rVert^{2}_{F}$. Let $\mathbf{U}\in\mathbb{R}^{m_{L}\times m_{l}}$ and $\mathbf{Q}\in\mathbb{R}^{m_{l}\times m_{l}}$ be matrices with elements sampled from $\mathcal{N}(0,1)$. We now fit $\hat{\mathbf{W}}_{1}$ to estimate $\mathbf{XQ}+\mathbf{YU}$ from $\mathbf{X}$, such that
$\hat{\mathbf{W}}_{1}:=(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}(\mathbf{XQ}+\mathbf{YU}).$ (38)
Thus, the pre-activation potential at the first layer is given by
$\mathbf{X}\hat{\mathbf{W}}_{1}=\mathbf{XQ}+\mathbf{P}_{\mathbf{X}}\mathbf{YU}.$ (39)
As $m_{l}\gg m_{L}$, we assume $\textbf{UU}^{+}=\textbf{I}_{m_{L}}$. Independence of $\mathbf{Q}$ and $\mathbf{U}$, with elements sampled from $\mathcal{N}(0,1)$ gives $\mathbb{E}_{\mathbf{Q},\mathbf{U}}\left[\mathbf{QU}^{+}\right]=0$. Therefore,
$\mathbb{E}_{\mathbf{Q},\mathbf{U}}\left[\mathbf{X}\hat{\mathbf{W}}_{1}\mathbf{U}^{+}\right]=\mathbb{E}_{\mathbf{Q},\mathbf{U}}\left[\mathbf{XQU}^{+}\right]+\mathbb{E}_{\mathbf{U}}\left[\mathbf{P}_{\mathbf{X}}\mathbf{YUU}^{+}\right]=\mathbf{P}_{\mathbf{X}}\mathbf{Y}.$ (40)
Therefore, we note that in the $m_{l}\gg m_{L}$ setting, $\mathbf{P}_{\mathbf{X}}\mathbf{Y}$ is encoded by random projection over $\mathbf{X}\hat{\mathbf{W}}_{1}$.
Let $\mathbf{A}=\operatorname{ReLU}(\mathbf{X}\hat{\mathbf{W}}_{1})$ denote the hidden activations. We note that $\operatorname{col}(\mathbf{A})\neq\operatorname{col}(\mathbf{X})$, as the $\operatorname{ReLU}$ operation introduces non-linearities into $\mathbf{X}\hat{\mathbf{W}}_{1}$. Since $(\mathbf{Q},\mathbf{U})$ is sampled from a symmetric distribution around zero, the realisations $(\mathbf{Q},\mathbf{U})$ and $(\mathbf{-Q},\mathbf{-U})$ have equal probability. Using the identity
$\mathbf{X}\hat{\mathbf{W}}_{1}=\operatorname{ReLU}(\mathbf{X}\hat{\mathbf{W}}_{1})-\operatorname{ReLU}(-\mathbf{X}\hat{\mathbf{W}}_{1}),$ (41)
we have
$\mathbb{E}_{\mathbf{Q},\mathbf{U}}\left[\operatorname{ReLU}(\mathbf{X}\hat{\mathbf{W}}_{1})\mathbf{U}^{+}\right]=-\mathbb{E}_{\mathbf{Q},\mathbf{U}}\left[\operatorname{ReLU}(-\mathbf{X}\hat{\mathbf{W}}_{1})\mathbf{U}^{+}\right].$ (42)
@@
$\mathbb{E}\left[\mathbf{A}\mathbf{U}^{+}\right]=\frac{1}{2}\mathbb{E}\left[\mathbf{X}\hat{\mathbf{W}}_{1}\mathbf{U}^{+}\right]=\frac{1}{2}\mathbf{P}_{\mathbf{X}}\mathbf{Y}.$ (43)
In the $m_{l}\gg m_{L}$ setting, we assume convergence to the mean, and approximate $\mathbf{A}\mathbf{U}^{+}\approx\frac{1}{2}\mathbf{P}_{\mathbf{X}}\mathbf{Y}$. Accordingly, we assume that $\mathbf{P}_{\mathbf{X}}\mathbf{Y}$ lies within $\operatorname{col}(\textbf{A})$ – i.e., $\operatorname{col}(\hat{\mathbf{Y}}_{\mathrm{linear}})\subset\operatorname{col}(\mathbf{A})$. We also assume $\mathbf{A}$ has full column rank. We now fit $\hat{\mathbf{W}}_{2}$ to estimate $\mathbf{Y}$ from $\mathbf{A}$, such that
$\hat{\mathbf{W}}_{2}:=(\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top}\mathbf{Y}.$ (44)
Letting $\hat{\mathbf{Y}}_{\mathrm{FP}}=\mathbf{P}_{\mathbf{A}}\mathbf{Y}$, the Forward Projection error is given by
$E_{\mathrm{FP}}=\mathbb{E}\left[\lVert\mathbf{Y}-\hat{\mathbf{Y}}_{\mathrm{FP}}\rVert^{2}_{F}\right]=\mathbb{E}\left[\lVert(\mathbf{I}-\mathbf{P}_{\mathbf{A}})\mathbf{Y}\rVert^{2}_{F}\right].$ (45)
From (35), we have
$(\mathbf{I}-\mathbf{P}_{\mathbf{A}})\mathbf{Y}=(\mathbf{I}-\mathbf{P}_{\mathbf{A}})\hat{\mathbf{Y}}_{\mathrm{linear}}+(\mathbf{I}-\mathbf{P}_{\mathbf{A}})\mathbf{Y}_{\perp}.$ (46)
From our assumption that $\operatorname{col}(\hat{\mathbf{Y}}_{\mathrm{linear}})\subset\operatorname{col}(\mathbf{A})$, we have
$(\mathbf{I}-\mathbf{P}_{\mathbf{A}})\hat{\mathbf{Y}}_{\mathrm{linear}}=0.$ (47)
Substituting (45) and (46), we have